BELL’S PRIMENESS CRITERION FOR W (2n +1)

MARK C. WILSON, GEOFFREY PRITCHARD, AND DAVID H. WOOD

Abstract. On the basis of experimental work involving matrix computations, we conjecture and prove that a criterion due to Bell for primeness of the universal enveloping algebra of a Lie superalgebra applies to the Cartan type Lie superalgebras W (n) for n = 3 but does not apply for odd n ≥ 5.